A reserach is observing cell division in cells derived from a diploid organism. She traps the cell at the end of mitotic prophase and counts a total of 108 chromatids

Based on this count, how many chromosomes would be contained in a normal gamete for this organism?

PS: the answer is not 54 I have already tried this answer.
Answer #1

The answer would be 27.

Since each chromosome would have two chromatids, 2 X Number of chromosomes = Total number of chromatids.

So, Number of chromosomes = 108/2 = 54.

In a gemete the number of chromosomes are reduced to half, hence the answer will be 54/2 = 27.

"A malaria biologist"